OpenStudy (anonymous):

solve for y, state the m and b 6x+2y=12

OpenStudy (anonymous):

First, you would set it up as "2y=6x+12"

OpenStudy (anonymous):

and divide all of them by 2

OpenStudy (anonymous):

point slope form for a line y = mx + b rewrite it in that form 6x - 12 = 2y 3x - 6 = y m = 3 b=6

OpenStudy (anonymous):

oops, thats -6

OpenStudy (anonymous):

m= the x and b = the other number that's not y.

OpenStudy (anonymous):

It's quite simple. You might want to learn this.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Just remember, it's always going to be Y= mx+b. that is the equation.